The current in an electric lamp is 5 A. What quantity of electricity flows towards the filament in 6 minutes?


1
Expert's answer
2021-10-04T04:55:17-0400

Q = I * t = 5 A * 6 * 60 sec = 1800 C

‘Quantity of electricity’ could be expressed as the number electrons. Each electron has charge 1.6E-19 C. Therefore Q = 1800/1.6E-19 = 1.125E+22 electrons = 1.125*10²² electrons.